Solve [2y^2 – 4x + 5]dx = [y – 2y^2 - 4xy]dy


1
Expert's answer
2021-05-17T12:04:02-0400

Given,

"[2y^2 \u2013 4x + 5]dx = [y \u2013 2y^2 - 4xy]dy"


Compare with "Mdx+Ndy=0"


"M=2y^2-4x+5,N=-(y^2-2^2-4xy)"


"\dfrac{dM}{dy}=4y, \dfrac{dN}{dx}=4y"


"\therefore \dfrac{dM}{dy}=\dfrac{dN}{dx}=4y"


The given equation is exact diff. equation. So Integrating all terms of M keeping y as a constsant and integrating only those part of N which are free from x.


So

"\int (2y^2-4x+5)dx-\int (y-2y^2-4xy)dy=0

\\[9pt]

2xy^2-2x^2+5x-\dfrac{y^2}{2}-\dfrac{2}{3}^3-2xy^2+C=0"


So Solution is-


"2xy^2-2x^2+5x-\dfrac{y^2}{2}-\dfrac{2}{3}^3-2xy^2+C=0"